OpenStudy (anonymous):

PLEASE HELP!! Solve 125x^3 - 27 We're doing cubics today and I know that first we have to factor them out and I got: (5x-3)(25x^2+15x+9) and we are supposed to get that trinomial. But then we have to solve it by either completing the square or the quadratic formula and I'm just getting stuck when I reach the square root part... Help would be much appreciated!! :) Thanks!!

OpenStudy (ksaimouli):



OpenStudy (anonymous):

@ksaimouli I can't do that lol

OpenStudy (anonymous):

this cubic function how to use completing the square or the quadratic formula ?????!

OpenStudy (anonymous):

are you trying to solve this EQUATION: \(\large 125x^3 - 27=0 \) ????

OpenStudy (anonymous):

@dpaInc Yep!

OpenStudy (anonymous):

ok... your factorization is good... \(\large 125x^3 - 27=0 \) \(\large (5x-3)(25x^2+15x+9)=0 \) solving for that first factor of (5x-3)=0 is trivial so i won't go over it.... unless you want me to... ??? but that second factor needs to be solved also: \(\large 25x^2+15x+9=0 \) using the quadratic formula...

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

@dpalnc Yeah I tried the quadratic formula and I got

OpenStudy (anonymous):